OpenStudy (anonymous):

A lottery winner must decide between two methods of payment. Choice one is to receive a lump sum of $100,000,000. Choice two is to begin with one cent on day 1, three cents on day 2, nine cents on day three, and so on until the end of 22 days. Which choice should the lottery winner select? Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

how should i go about to solve this?

OpenStudy (anonymous):

I guess that first you should figure out how much would she gain by receiving one cent on day 1, three cents on day 2, and so on.

OpenStudy (anonymous):

3^22 = ?

OpenStudy (anonymous):

Actually, 3^21 I guess because day 1 = 1 c.

OpenStudy (anonymous):

I guess you should use the sum formula of sequencies

OpenStudy (anonymous):

i kno that the lesson is talking about geometric sequences so i have to explain it in that way but idk how it relates o.O

OpenStudy (anonymous):

yeah, exactly

OpenStudy (anonymous):

do you know the GS sum formula?

OpenStudy (anonymous):

so i would use the formula an=a1 (r)^n-1 or sn= a1-anr/1-r?

OpenStudy (anonymous):

sn

OpenStudy (anonymous):

so you can know how much she would gain in those 22 days

OpenStudy (anonymous):

3^21 (just last days payment) is so much bigger than 100,000,000. So, why bother doing Geometric series?

OpenStudy (anonymous):

I don't know if that's right, GT. Besides, the exercise wants to test her ability to apply the formula, which can seem not important but it actually is. There are some exercises that she will only get them using the formulas.

OpenStudy (anonymous):

i have to use geometric series because that wat the teach wants :P

OpenStudy (anonymous):

lol that too^

OpenStudy (anonymous):

OK. Plug in the formula.

OpenStudy (anonymous):

K. I need to go now, keep helping her! Bye guys, and good luck

OpenStudy (anonymous):

kk thanks bye

OpenStudy (anonymous):

3^21 = 10,460,353,203